Orange light with a wavelength of 590 nm passes through a single slit creates a pattern on a screen located some distance away so that the first dark fringe occurs at an angle of 8.0o What is the width of the slit? (


1
Expert's answer
2021-04-12T07:03:58-0400

The equation for the minimums of the difraction on one slit is the following (see https://courses.lumenlearning.com/physics/chapter/27-5-single-slit-diffraction/):


"D\sin \theta = m\lambda"

where "D" is the width of the slit, "m = 1" is the order of minimum (first dark fringe), "\theta = 8\degree" is the angle of this minimum, and "\lambda = 590nm = 5.9\times 10^{-7}m" is the wavelength of the light. Thus, obtain:


"D = \dfrac{m\lambda}{\sin\theta}\\
D = \dfrac{1\cdot 5.9\times 10^{-7}}{\sin8\degree} \approx 42\times 10^{-7}m = 4.2\mu m"

Answer. "4.2\mu m".